Calculate the Molarity of each of the following solutions:


a. 185 g of sucrose, C12H22011 in 1.00 L of solution


b. 6.30 g of HNO3 dissolved in 255 ml of solution.

1
Expert's answer
2022-03-03T04:45:04-0500

a) moles= mass/ rmm= 185/342= 0.54 moles

Molarity = moles/ litre

= 0.54/1= 0.54M

b)moles= 6.3/63= 0.1 moles

Molarity = 0.1/0.255= 0.39M